OpenStudy (yatonoragami):

Circle A has a diameter of 7 inches, a circumference of 21.98 inches, and an area of 38.465 square inches. The diameter of circle B is 6 inches, the circumference is 18.84 inches, and the area is 28.26 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

OpenStudy (welshfella):

C = pi * d 21.98 = pi * 7

OpenStudy (welshfella):

- solve for pi Do similar thing for circle B.

OpenStudy (yatonoragami):

So would I do 21.98 x 3.14 x 7?

OpenStudy (yatonoragami):

And what do you mean with C= pi * d?

OpenStudy (welshfella):

No you have to isolate pi do this by dividing both sides of the equation by 7 

OpenStudy (yatonoragami):

So I would do 21.98 divided by 7??

OpenStudy (welshfella):

C = pi * d means Circumference = pi times diameter

OpenStudy (welshfella):

yes

OpenStudy (yatonoragami):

That would be pi?